OpenStudy (anonymous):

How do you find an answer such as yes or no to, 84/112 and 168/192 being equivalent?

OpenStudy (anonymous):

84/112 = 3/4 168/192 = 7/8 since 3/4 =/= 7/8 they aren't equivalent

OpenStudy (anonymous):

I know, but how do you find it? I simplified it to 3/4 but how do you know if the number is higher?

OpenStudy (anonymous):

fractions

OpenStudy (anonymous):

Do you know how to factor a number like 84? (write it as the product of prime numbers)